How do I solve for x in this problem?
1/x + 1/(x - 3) = 2/(x + 3)

Answer by nerdybill(7384)

1/x + 1/(x - 3) = 2/(x + 3)
Multiply both sides by x(x-3)(x+3) to get:
(x-3)(x+3) + x(x+3) = 2x(x-3)
(x^2-3x+3x-9) + (x^2+3x) = (2x^2-6x)
x^2-3x+3x-9 + x^2+3x = 2x^2-6x
2x^2+3x-9 = 2x^2-6x
3x-9 = -6x
9x-9 = 0
9x = 9
x = 1